OpenStudy (anonymous):

What are the solutions of the quadratic equation? x2 – 9x + 20 = 0 ?

hartnn (hartnn):

for factorizing the standard quadratic equation,\( ax^2+bx+c=0\), we need 2 numbers whose sum is 'b' and product = ac for your Question, can you find 2 numbers whose sum = -9 and product = 20 ?

OpenStudy (anonymous):

-5 and -4 ?

hartnn (hartnn):

correct! so split up -9x as -5x-4x x^2-5x-4x+20 =0 can you factor out 'x' from 1st two terms ?

OpenStudy (anonymous):

I don't know how

OpenStudy (anonymous):

would I do 5+(-1x)=0 ?

hartnn (hartnn):

x^2-5x-4x+20 =0 x (x-5) -4 (x-5) = 0 see what i did ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

now factor out x-5 from that ? what do u get ?

OpenStudy (anonymous):

x-4=0 ?

hartnn (hartnn):

x (x-5) -4 (x-5) = 0 (x-5)(x-4) =0 so, x-5 = 0 x-4 = 0 x=5, 4

OpenStudy (anonymous):

Thank you ( :

hartnn (hartnn):

welcome ^_^